# Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

### v1

@@ 3.3 Executor
Each robot is launched with its own individual executor. As shown in Fig. 3, each executor consists of an RPC client, an action queue, and a controller. The RPC client establishes communication with the server, receiving enqueued actions and sending confirmation for finished actions. The action queue keeps track of the enqueued actions, which are executed in a first-in-first-out manner. A Proportional Integral Derivative (PID) controller is used to execute these actions by sending control commands to the simulator. Once an action is finished, a confirmation message is sent to the server. To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue.
3.4 Performance Metrics
@@
SMART provides several built-in performance metrics, including (1) Average execution time (AET): The sum of the simulation times required by all robots to complete their assigned paths, divided by the number of robots, (2) Maximum execution time: The longest simulation time taken by any single robot to complete its path, and (3) Robot state over time: The state of an individual robot at different times, where the state of a robot includes the size of its action queue, its current position, and control output. Additionally, the open-source nature of our code allows researchers to easily derive custom metrics from SMART as needed. SMART also provides visualization for the MAPF plan executions, allowing researchers to better understand the behaviors of the robots.
4 System Evaluation
@@
We test SMART on two machines: (1) a server with a 64-core AMD 7980X processor and 256 GB of memory, and (2) a laptop with a 4-core Intel i7-6700HQ processor and 16 GB of memory. Both machines are equipped with Ubuntu 20.04. We use machine (1) to evaluate the scalability and applicability of SMART in Sections 4.1 and 4.2. We use both machines to evaluate CPU and memory usage in Section 4.3.
4.1 Scalability and Runtime Performance
This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the total number of simulation updates to the real-world time (in seconds). A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands. In all experiments, the simulation update period is set to 0.1 simulation seconds. A higher simulation speed shows that more simulation updates can be finished within the same unit of time, indicating better scalability. For example, if the simulation speed is 100 and the update period is 0.1 simulation seconds, the simulation world runs 10 times faster than real time. We assess scalability on two environments converted from: warehouse (warehouse-10-20-10-2-1, size: 161$\times$63) from the MovingAI benchmark Stern et al. (2019) and sortation-center (size: $500\times 140$) from the League of Robot Runner MAPF competition Chan et al. (2024). MAPF plans are generated using MAPF-LNS2222Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 Li et al. (2022), a state-of-the-art MAPF planner, with a progressively increasing number of robots and a time limit of 60 seconds. These plans are then executed in SMART without visualization. Notably, the MAPF instances from warehouse support a maximum of 1,000 robots, whereas MAPF-LNS2 can handle up to 2,000 robots within the time limit on sortation-center.
@@ 4.1 Scalability and Runtime Performance
Figure 5: Simulation speed on warehouse and sortation-center maps.
As shown in Fig. 5, SMART supports simulations with up to 2,000 robots. The more robots and the larger the map, the smaller the simulation speed. Nevertheless, SMART maintains a simulation speed of around 100 with 1,000 robots in warehouse. Even in the larger sortation-center, the simulation speed is consistently greater than 10.
Robots (#) average execution time maximum execution time
@@
Table 1: Replicability using different numbers of robots.
4.2 Replicability
@@
We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs. Due to the realistic setting we used, uncertainty may arise from factors such as network communication delays or controller variability. Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan. For each run, we record both the average execution time and the maximum execution time. As shown in Table 1, SMART demonstrates strong replicability across runs, with minor execution time variations. While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
4.3 CPU and Memory Usage
@@
We evaluate the CPU and memory usage of SMART on both machines (1) and (2). We use machine (2) to demonstrate that SMART can run on resource-constrained machines. To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark Stern et al. (2019). All simulations are parallelized on all CPU cores on both machines. We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots. Table 2 shows the results. We observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops. In machine (2), the simulation speed is slower than in machine (1) due to an inferior processor. For reference, we run the Gazebo-based simulator Varambally et al. (2022) on machine (2) with 50 robots in an environment of size 31$\times$12. Since Gazebo operates on a single thread, it utilizes 100% of a single CPU core, consumes 2.09 GB of memory, and achieves a simulation speed of 5.40.
Machine Robots (#) CPU Usage (%) Memory Usage (GB) Simulation Speed
(1) 100 16.48 $\pm$ 0.58 0.19 $\pm$ 0.26 334.90 $\pm$ 1.54
1000 46.02 $\pm$ 0.98 0.65 $\pm$ 0.03 73.71 $\pm$ 0.40
(2) 100 27.96 $\pm$ 0.39 0.04 $\pm$ 0.00 143.74 $\pm$ 1.03
1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 19.45 $\pm$ 0.09
Table 2: Comparison of CPU and Memory Usage
5 Problems from MAPF to Realistic Scenarios
We argue that SMART has several broader impacts on the MAPF research community. First, it allows researchers to evaluate MAPF plans from arbitrary MAPF planners in a simulator with realistic robot kinodynamics models and execution delays caused by real-world factors. Second, it bridges the gap between MAPF planning algorithms and executing the planned paths in the real world, bringing researchers more guidance on designing high-quality MAPF algorithms for the real world. Third, SMART can be used to study research problems regarding executing MAPF plans in the real world.
In this section, we study research problems focusing on the gap between MAPF planning and real-world execution of the planned paths. We study these problems using six simulation environments converted from six maps from the MovingAI benchmark Stern et al. (2019): empty (empty-32-32, size: 32$\times$32), random (random-64-64-10, size: 64$\times$64), room (room-64-64-16, size: 64$\times$64), den312d (den312d, size: 65$\times$81), maze (maze-32-32-4, size: 32$\times$32), and warehouse (warehouse-10-20-10-2-1, size: 161$\times$63).
Feature All Features SOC Rotations (#) Type-1 edges (#) Type-2 edges (#) Conflict robot pairs (#) Robots (#)
MAPE 0.0342 0.1778 0.1793 0.2414 0.2826 0.3175 0.4317
Table 3: Quadratic Regression Results. Except for the first column, each column means that the model is trained with only a single feature
5.1 Impact of Optimality
We first use SMART to study the practical effectiveness of the objective used by MAPF algorithms. One of the most widely used objectives in MAPF research is the Sum of Costs (SOC), which measures the total number of timesteps for all robots to reach their goals. While extensive efforts have been made to develop algorithms that optimize SOC optimally or sub-optimally, it is unclear, for example, whether reducing SOC by 1% translates to a similar degree of improvement in AET in a more realistic simulation. In this section, we use SMART to assess the correlation between improvements in SOC and AET. Moreover, we want to explore better objectives than SOC in terms of evaluating the quality of the MAPF plan in more realistic settings.
5.1.1 Experiment Setup
We generate MAPF plans with varying SOC for a specific MAPF instance by using MAPF-LNS [^3] Code at https://github.com/Jiaoyang-Li/MAPF-LNS Li et al. (2021a), an anytime method that iteratively refines MAPF plans to improve SOC within a predefined runtime limit. For each map, we select three different numbers of robots and five sets of different start and goal locations for robots. For each MAPF instance, we record all plans generated in each iteration of MAPF-LNS within the runtime limit of 60 seconds. We then execute these plans in SMART.
Figure 6: Relationship between SOC and AET. Each dot represents a MAPF plan and its execution result.
5.1.2 Results and Analysis
We first study whether SOC can measure the real-world execution time of a MAPF plan. As shown in Fig. 6, SOC and AET show a strong positive correlation across all environments and all numbers of robots, indicating that SOC captures the trend of AET. However, the correlation is not perfectly monotonic. For the same SOC value, there are noticeable variations in AET, suggesting that SOC alone cannot account for all factors influencing execution time. This highlights the limitation of SOC as the sole objective of the MAPF problem.
To explore other factors that influence execution time, we extract additional features from the ADG, including (1) Type-2 edges (#), the number of type-2 edges, (2) Type-1 edges (#), the number of type-1 edges, (3) Rotations (#), the number of rotation actions, (4) Conflict robot pairs (#), the number of robot pairs connected by type-2 edges, reflecting dependencies in their passing orders, and (5) Robots (#), the number of robots. We analyze the correlation between these features and AET by using a quadratic regression model to predict AET based on these features. We measure prediction accuracy by using the mean absolute percentage error (MAPE). As shown in Table 3, SOC exhibits the highest predictive power among individual features, while other features, such as the number of Type-2 edges, Type-1 edges, and Rotations, also show correlation with AET. Meanwhile, the model incorporating all features achieves a much lower MAPE than any single feature, indicating that multiple features influence the execution time. In conclusion, incorporating features in addition to SOC is crucial to developing a more accurate objective function for execution time estimation during MAPF planning.
Figure 7: AET comparison for MAPF with different robot models.
5.2 Impact of Different Robot Models
The standard MAPF problem relies on a simplified model that ignores critical execution factors such as kinodynamics and potential delays. This simplification creates a gap between planning and real-world execution. Thus, recent research has introduced more realistic MAPF models that consider real-world factors. An ideal MAPF model should capture the essential aspects of robot kinodynamics while omitting unnecessary details. As SMART supports various robot models, we conduct an empirical comparison of different MAPF models to identify the most effective one in real-world scenarios. In Varambally et al. (2022), a similar problem is studied, but the experiments are limited to a warehouse map and consider no more than 50 robots. Additionally, we compare a wider range of robot models.
5.2.1 Experiment Setup
We evaluate the impact of incorporating robot kinodynamics and potential delays. For the kinodynamics models, we consider: (1) the standard MAPF model, (2) the MAPF model with rotation, and (3) the MAPF model that accounts for the full kinodynamics of robots, including rotation, speed, and acceleration. To model delays, we use the $k$-robust delay model Atzmon et al. (2020), which ensures that robots can avoid collisions even if delayed by up to $k$ timesteps. By combining kinodynamics and delay factors, we evaluate the following MAPF algorithms: (1) MAPF, $k=i$: This method uses the standard MAPF model and a $k$-robust delay model of $k=i$. (2) MAPF+rotation, $k=i$: This extends the standard MAPF model by incorporating in-place rotation but does not account for other kinematics. It also includes a $k$-robust delay model. (3) MAPF+kinodynamics: Derived from Yan and Li (2024a), this method incorporates accurate kinodynamic constraints of robots. We use PBS Ma et al. (2019) to handle conflicts between robots under these models with a runtime limit of 60 seconds. We run these methods to generate MAPF plans for all maps using 25 “random scenarios” with a progressive number of robots until the success rate is smaller than 20%. Finally, these plans are executed using SMART.
Figure 8: Paths generated on different maps.
5.2.2 Results and Analysis
As shown in Fig. 7, the method incorporating the full kinodynamics of robots achieves the shortest AET among all the models. Additionally, the planner considering robot rotations demonstrates much better AET than the standard MAPF model. These results indicate that incorporating more accurate kinodynamic models into planning can improve execution time. However, we also observe trade-offs associated with these models. More complex kinodynamic models lead to longer runtimes during planning. Meanwhile, the $k$-robust model achieves better or comparable performance with fewer robots with larger values of $k$; however, its performance declines significantly as the number of robots increases, resulting in lower solution quality and reduced scalability. Interestingly, in warehouse environments, MAPF+kinodynamics does not provide significant improvements over the MAPF+rotation. To better understand this, we visualize the MAPF plans in Fig. 8, comparing the warehouse map with random and room maps. The planned paths show that MAPF+rotation and MAPF+kinodynamics generate paths with more consecutive movements and fewer conflict vertices in warehouse maps than in the other two maps. We hypothesize that this minimizes the potential benefits of accounting for kinodynamics in such environments. In summary, while more accurate kinodynamic models generally improve execution time in realistic settings, their increased computational complexity can limit scalability. Additionally, the $k$-robust model is less efficient and may even harm performance, particularly in more congested scenarios.
Figure 9: AET using different MAPF models and planners.
5.3 Optimality and Robot Model
From Sections 5.1 and 5.2, we show that both improving SOC and incorporating a more accurate kinodynamic model during planning can lead to better AET during execution. However, both of them require additional computation time. Considering the limited planning time available in real-world applications, this section explores the balance between these factors to provide practical insights.
5.3.1 Experiment Setup
We generate MAPF plans with varying SOC for three robot models: (1) standard MAPF, (2) MAPF with rotation, and (3) MAPF with kinodynamics. To obtain optimal solutions, we use CBS [^4] Code at https://github.com/Jiaoyang-Li/CBSH2-RTC Li et al. (2021b) for the standard MAPF model and CBS with rotation [^5] Code at https://github.com/YueZhang-studyuse/MAPF_T Zhang et al. (2023a) for the rotation model. For suboptimal plans, we combine PBS with all three models, as described in Section 5.2. To analyze the relationship between AET and SOC, we create a naive anytime planner by running PP with random restarts and recording the MAPF plan with the lowest SOC until reaching a runtime limit of 60 seconds. We do not use existing anytime solvers, such as MAPF-LNS, because, to our best knowledge, no available implementation considers robot models with kinodynamic constraints.
5.3.2 Results and Analysis
As shown in Fig. 9, planners using more accurate robot models consistently achieve better AET across all maps, even when the methods using less accurate models are optimal. This indicates that incorporating more accurate kinodynamic robot models into planning may be more important than obtaining optimal solutions with less accurate models. Notably, as discussed in Section 5.1, since SOC is not a perfect metric, the optimal planner does not always achieve the best AET. Meanwhile, execution frameworks may also become a bottleneck. Since ADG adopts an overly conservative strategy, assuming robots can be delayed indefinitely, it compromises the solution quality of the MAPF plan in certain cases. In conclusion, the limiting factors for achieving better AET lie in the more accurate robot models or the execution frameworks with better performance.
6 Conclusion
@@
We introduced SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings. SMART bridges the gap between methods that use simplified models and their real-world deployments. SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments. Our experiments demonstrated that SMART can handle thousands of robots while maintaining high simulation speeds. We highlight several key challenges arising from the application of MAPF in realistic scenarios. We further show how SMART can be used to address these challenges. Future work includes extending the framework to address lifelong MAPF scenarios, integrating it with other execution frameworks Coskun and O’Kane (2019); Berndt et al. (2023); Feng et al. (2024), and interfacing it with real robots.
Acknowledgments
The research was supported by the National Science Foundation (NSF) under grant number #2328671 and a gift from Amazon. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.

### v2

@@ III-C Executor
Each robot is launched with an individual executor. As shown in Fig. 3, each executor consists of an RPC client, an action queue, and a controller. The RPC client establishes communication with the server, receiving enqueued actions and sending confirmation for finished actions. The action queue keeps track of the enqueued actions, which are executed in a first-in-first-out manner. A Proportional Integral Derivative (PID) controller is used to execute these actions by sending control commands to the simulator. Once an action is finished, a confirmation message is sent to the server. To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue. For example, if a robot receives three consecutive forward movement actions, the executor merges them into a single longer forward action, reducing unnecessary stop-and-go behavior.
III-D Performance Metrics
@@
SMART provides several built-in performance metrics, including (1) Average execution time (AET): The sum of the simulation times required by all robots to complete their assigned paths, divided by the number of robots, (2) Maximum execution time: The longest simulation time taken by any single robot to complete its path, and (3) Robot state over time: The state of each robot at different times, where the state includes the size of its action queue, its current position, and its current wheel velocity. Additionally, the open-source nature of our code allows researchers to easily derive custom metrics from SMART as needed. SMART also provides visualization for the MAPF plan executions, allowing researchers to better understand the behaviors of the robots.
IV Using SMART
Users can evaluate their solutions using SMART in two ways: by uploading their MAPF problem and solution through a web interface or by deploying SMART locally.
Figure 5: The SMART web interface visualizing a multi-robot simulation on a grid map. In the viewport, users interact with individual robots to inspect paths and inter-robot constraints. On the side panel, users can configure the environment, control playback, and see execution details.
SMART as a Service
SMART, with ARGoS3 and the 3D visualizer, is directly available as software-as-a-service [^2] The link will be made available here upon acceptance.. Specifically, we provide an intuitive web app for running and inspecting MAPF simulations. Complete with a beautiful 3D-visualizer built with React and Three.js, it allows researchers to instantly understand how algorithms perform with physical robots (Fig. 5). Users configure the environment by specifying a map and scenario file in the MovingAI Benchmark format. Within the online interface, after users upload the files and select the desired kinodynamic settings for the robots, and initiate the simulation with a single click. After starting the simulation, they can control the simulation time step and adjust the playback speed. Users may select robots to view complete paths, execution status, and Type-2 edges between robots.
Installing SMART Locally
Since SMART provides flexible interfaces for customization, researchers may choose to compile it locally for advanced use cases, such as running it in headless mode, integrating it into custom pipelines, or interfacing with real robots. As an open-source tool, SMART includes documentation and examples to support these advanced configurations.
V System Evaluation
@@
We test SMART on two machines. All experiments are conducted on machine (1), a high-performance server with a 64-core AMD 7980X processor, 256 GB of memory, and RTX 3090Ti GPU. In Section V-C, we additionally test SMART on machine (2), a low-end laptop with a 4-core Intel i7-6700HQ processor and 16 GB of memory, to demonstrate SMART’s performance on resource-constrained hardware. Both machines are equipped with Ubuntu 20.04.
@@
Figure 6: Simulation environments. (a) empty (b) maze (c)random (d) game (e) warehouse (f) sortation-center.
[…]
In all experiments, the simulation update period is set to 0.1 simulation seconds for both simulators. A simulation update refers to the periodic recalculation of the status of robots and the environment based on the input commands. In ARGoS3, each grid cell represents 1 meter. Each foot-bot robot is subject to a speed limit of [0, 5] m/s, an acceleration limit of [-0.4, 0.4] m/s2, and an angular velocity limit of 30/s. In Isaac Sim, each grid cell represents 0.5 meters. Each JetBot is constrained to a speed limit of [0, 2.5] m/s, an acceleration limit of [-0.2, 0.2] m/s2, and the same angular velocity limit of 30/s.
We evaluated our system using MAPF plans generated by three different methods: CBS [12], PBS [7], and MAPF-LNS2 [9]. These plans were created using four different robot models: the standard MAPF model [1], the $k$-robust delay model [19], MAPF with rotation [33], and MAPF with kinodynamics [21]. Across all combinations, our system achieved a 100% execution success rate. In the remainder of this section, we present results based on MAPF plans generated by MAPF-LNS2, as it demonstrated the highest scalability among the algorithms evaluated.
V-A Scalability and Runtime Performance
This section evaluates the scalability and runtime performance of SMART by analyzing its simulation speed, which is defined as the ratio of the simulation time to the real-world time. A higher simulation speed means that more simulation updates can be finished within the same unit of time, indicating better scalability. For example, if the simulation speed is 10, the simulation world runs 10 times faster than in real time. MAPF plans are generated using MAPF-LNS2333Code at https://github.com/Jiaoyang-Li/MAPF-LNS2 [9] with a progressively increasing number of robots and a time limit of 60 seconds. For each map, we randomly select 5 scenarios from the MovingAI Benchmark to generate plans, which are then executed in SMART without visualization. In ARGoS3, the maximum number of robots tested corresponds to the limit imposed by the MAPF instances from the MovingAI Benchmark for all maps except sortation-center, which is set to 2,000 robots. In contrast, for Isaac Sim, the upper bound on the number of robots is determined by the point at which the simulation speed drops below real-time (i.e., simulation speed << 1).
Simulator Robots average execution maximum execution
(#) time time
ARGoS3 10 123.33 $\pm$ 0.00 373.6 $\pm$ 0.00
50 182.84 $\pm$ 0.01 433.30 $\pm$ 0.00
@@ V-A Scalability and Runtime Performance
Isaac Sim 10 122.47 $\pm$ 0.00 369.47 $\pm$ 0.00
50 183.71 $\pm$ 0.07 435.46 $\pm$ 1.36
TABLE I: Replicability using different numbers of robots.
Figure 7: Simulation speed on all environments.
As shown in Fig. 7, SMART supports simulations with up to 2,000 robots. The more robots and the larger the map, the smaller the simulation speed. Nevertheless, for ARGoS3, SMART maintains a simulation speed of around 10.0 with 1,000 robots in warehouse. Even in the larger sortation-center, the simulation speed is consistently greater than 1.0. In contrast, Isaac Sim provides the capability to model more complex real-world factors while also producing photorealistic visualizations, but it is significantly less scalable, particularly in large environments. It is therefore more suitable for experiments with a smaller number of agents.
V-B Replicability
@@
We evaluate the consistency of SMART by examining whether the same MAPF plans yield the same outcomes across multiple runs. Due to the realistic setting we used, uncertainty may arise from factors such as network communication delays or controller variability. Using the same warehouse map and MAPF plans as in the previous section, we execute SMART 10 times for each MAPF plan. For each run, we record both the average and the maximum execution time. As shown in Table I, SMART demonstrates strong replicability across runs, with minor execution time variations. While these variations increase slightly with more robots, they remain minimal even for up to a thousand robots.
V-C CPU and Memory Usage
@@
We evaluate the CPU and memory usage of SMART on both machines (1) and (2). We use machine (2) to demonstrate that SMART can run on resource-constrained machines. To evaluate usage, we generate MAPF plans with MAPF-LNS2 and run SMART on the warehouse map with different numbers of robots, each with 5 random MAPF instances selected from the MAPF benchmark [1]. All simulations are parallelized on all CPU cores on both machines. We record the peak CPU and memory usage while running the simulation and compute the average of all instances for each number of robots. Table II shows the results. When using ARGoS3 as the simulator, we observe that SMART has reasonable CPU usage and low memory usage on both machines, making it possible to run on low-spec personal laptops. The simulation speed in machine (2) is slower than in machine (1) due to an inferior processor. In contrast, simulations with Isaac Sim result in significantly higher CPU and memory usage, indicating that Isaac Sim is better suited for scenarios with fewer agents but more powerful hardware.
Sim Robots CPU Usage Memory Usage Simulation Speed
(#) (%) (GB)
[…]
100 16.48 $\pm$ 0.58 0.19 $\pm$ 0.26 33.49 $\pm$ 0.15
500 22.86 $\pm$ 1.59 0.24 $\pm$ 0.01 13.21 $\pm$ 0.13
1000 46.02 $\pm$ 0.98 0.65 $\pm$ 0.03 7.37 $\pm$ 0.04
[…]
100 N/A N/A N/A
500 N/A N/A N/A
1000 N/A N/A N/A
[…]
50 20.20 $\pm$ 0.91 0.05 $\pm$ 0.03 21.42 $\pm$ 0.35
100 27.96 $\pm$ 0.39 0.04 $\pm$ 0.00 14.37 $\pm$ 0.10
500 68.74 $\pm$ 17.24 0.22 $\pm$ 0.01 3.89 $\pm$ 0.02
1000 59.64 $\pm$ 0.32 0.63 $\pm$ 0.02 1.95 $\pm$ 0.01
TABLE II: Comparison of CPU usage, memory consumption, and simulation speed for ARGoS3 and Isaac Sim across two machines and various robot counts. Since Isaac Sim only supports recent GPUs with ray tracing capabilities, we conducted experiments with Isaac Sim only on Machine (1).
Figure 8: (a) Experimental setup in Isaac Sim. (b) Experimental setup with real robots.
V-D Evaluation in Real World
[…]
Given a set of start and goal locations, MAPF-LNS2 is used on a central server to generate a MAPF plan. SMART then processes this plan and transmits control commands to each robot via a wireless network, with a control frequency of 40 Hz. Each JetBot follows its assigned commands accordingly.
As shown in the accompanying multimedia material, SMART robustly executes the MAPF plan on physical robots, highlighting the practical feasibility and reliability of our planning framework in real-world scenarios. Across 10 trials, SMART achieved a 100% success rate, with an average execution time of 48.48 $\pm$ 8.79 seconds and a maximum execution time of 73.83 $\pm$ 9.81 seconds. For reference, in Isaac Sim, the average execution time is 45.66 $\pm$ 0.01 seconds, and the maximum execution time is 70.97 $\pm$ 0.00 seconds.
VI Conclusion
@@
We introduce SMART, a scalable testbed for evaluating MAPF algorithms under realistic settings. SMART bridges the gap between methods that use simplified MAPF models and their real-world deployments. SMART is compatible with existing 2-D grid-based benchmarks, making it straightforward for researchers to evaluate their methods in more realistic environments. We integrate SMART with two simulators, ARGoS3 and Isaac Sim. Our experiments show that SMART can handle thousands of robots while maintaining high simulation speeds when using ARGoS3. Future work includes extending the framework to address lifelong MAPF scenarios and integrating it with other execution frameworks [34, 35, 36].